OpenStudy (anonymous):

A body is moving in simple harmonic motion. Its position is modeled by s=sin(t)+cos(t), where s is measured in meters and t in seconds a. Find the body’s velocity, speed, acceleration, and jerk at time t=pi/4 seconds b. Find the time when the velocity is zero on [0,5] c. Describe the motion of the body from t = 0 to t = 5 seconds

OpenStudy (phi):

did you find the first and 2nd derivatives?

OpenStudy (phi):

and 3rd derivative of s for "jerk"

OpenStudy (anonymous):

PART A velocity(time)= s'(t)=v(t)=cos(t)-sin(t) a(t)=v'(t)=s''(t)=-sin(t)-cos(t) j(t)=a'(t)=s'''(t)=-cos(t)+sin(t) v(pi/4)=cos(pi/4)-sin(pi/4)=[sqrt(2)/2]-[sqrt(2)/2]=0 a(pi/4)=-sin(pi/4)-cos(pi/4)=-[sqrt(2)/2]-[sqrt(2)/2]=-sqrt(2) j(pi/4)=-cos(pi/4)+sin(pi/4)=-[sqrt(2)/2]+[sqrt(2)/2]=0

OpenStudy (anonymous):

ok i have to go now but ill be on so look later! thanks:)

OpenStudy (anonymous):

*to

OpenStudy (anonymous):

PART B v(t)=0=cos(t)-sin(t) cos(t)=sin(t) tan(t)=1 t={pi/4; pi+pi/4}